OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? i think its D but im not sure about B and D because i got close to the answer. A. 3, 4, 5 B. 28, 26, 12 C. 17, 14, 6 D. all of the above

OpenStudy (anonymous):

A. 3, 4, 5 B. 28, 26, 12 C. 17, 14, 6 D. all of the above

OpenStudy (anonymous):

i meant, im not sure about B and C not D i think its D.

OpenStudy (phi):

You use pythagorean theorem and a calculator. Does 12*12 + 26*26= 28*28? if we type 12*12 + 26*26= in the google search window (it does calculations) we get 820 if we type in 28*28= we get 784. 820 is not 784, so we do not have a right triangle. on the other hand: 3*3+4*4 = 5*5? 9+16= 25 5*5= 25 so 3,4,5 is a right triangle